OpenStudy (anonymous):

A box shaped like a rectangular prism has a height of 17 in and a volume of 2,720 in³ The length is 4 inches greater than twice the width. What is the width of the box?

OpenStudy (radar):

Volume V= length(l) X width (w) X height (h) V=lwh 2720 cu in. =17 X w X (2w+4) 2720 =17w(2w+4)=34w^2+68w=2760

OpenStudy (anonymous):

...? r u sure, the answers I got are a lot different, –10 in 8 in 20 in 40 in One of those... *confusedd*

OpenStudy (radar):

Typo change the 2760 to 2720

OpenStudy (anonymous):

Okay, okay, lol, thanks!!

OpenStudy (anonymous):

It's 8!!! :) I think.. lol

OpenStudy (radar):

\[34w ^{2}+68w-2720=0\]You want to solve for w. Divide the equation by 34 getting:\[w ^{2}+2w-80=0\] This factors quite nicely: (w+10)(w-8)=0 w=8 as negative value for w would be inappropriate. w=8 l=20

OpenStudy (radar):

Yes you beat me to it, great and good luck with your studies.

OpenStudy (anonymous):

Thanks you!!:)